# Jiwen Yu1∗†  Yiran Qin1∗   Xintao Wang2‡   Pengfei Wan2  Di Zhang2  Xihui Liu1‡

### v3

@@ 3 Preliminaries
$\mathcal{L}_{\mathbf{a}}(\boldsymbol{\phi})=\mathbb{E}[||\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{Z}_{t},\mathbf{p},\mathbf{A},t)-\boldsymbol{\epsilon}||_{2}^{2}],$ (1)
@@ 4.2 Action Control Module
[…]
❑ Grouping Actions with a Sliding Window. Due to the temporal compression ratio $r$, the number of actions ($rn$) differs from the number of features ($n+1$), creating a granularity mismatch for action-feature fusion. As shown in Fig. 4, we address this by grouping actions using a sliding window of size $w$. For the $i$-th feature $\mathbf{f}^{i}$, we consider actions within $[\mathbf{a}^{r\times(i-w+1)},...,\mathbf{a}^{ri}]$. This window design captures delayed action effects, such as how a jump command influences multiple subsequent frames. For out-of-range indices, boundary actions are used as padding. For mouse movement $\mathbf{M}$, the grouped action is $\mathbf{M}_{group}\in\mathbb{R}^{(n+1)\times rw\times d_{1}}$. As for keyboard actions $\mathbf{K}$, we first learn the embedding of actions and add the positional encoding. After that, we perform a grouping operation on the action embeddings to get $\mathbf{K}_{group}\in\mathbb{R}^{(n+1)\times rw\times c}$.
@@ 4.3 Autoregressive Generation of Long Action-Controllable Game Videos
Figure 5: Illustration of autoregressive video generation. The frames from index 0 to $k$ serve as conditional frames, while the remaining $N-k$ frames are for prediction, with $k$ randomly selected. (a) Training stage: Loss computation and optimization focus only on the noise of predicted frames. (b) Inference stage: The model iteratively selects the latest $k+1$ frames as conditions to generate $N-k$ new frames, enabling autoregressive generation.
@@
As shown in Fig. 5 (a), during training with $N+1$ frame latents (from 0 to $N$), we randomly select the first $k+1$ frames as conditions without adding noise, while adding noise only to the remaining $N-k$ frames for noise prediction training. Although the first $k+1$ frames are input to the model, since they are assumed to be already generated, their predicted noise outputs are not utilized. To improve training efficiency, we focus on computing training losses only for the $N-k$ frames that require prediction. As for inference, as shown in Fig. 5 (b), after full-sequence generation of the first $N+1$ frame latents, we can autoregressively generate new $N-k$ frames by selecting the most recent $k+1$ frames from history video latents as conditions for each subsequent generation, and merge them into the history video latents. This process can be repeated to achieve infinite-length video generation. Unlike conventional next-frame generation methods [8, 42, 11], our approach supports multi-frame generation in one step, greatly reducing the time for long video generation.
@@ 5.2 Multi-Phase Training Strategy
Resolution 280×150280$\times$ 150280 × 150 240240240p 720720720p 640×360640$\times$ 360640 × 360 720720720p 720720720p 640×360640$\times$ 360640 × 360
@@ 6.1 Implementation Details
❑ Pretrained Model Setting. Our experiments are based on an internal  [^1] B-sized transformer-based text-to-video diffusion model for research purpose, which is distilled from a larger pretrained video diffusion model and possesses strong generative priors in the open domain. The resolution of the game videos is 360×640360$\times$ 640360 × 640. The temporal compression rate of the VAE is $r=4$.
❑ Training and Inference Setting. Each phase of fine-tuning or training requires about two-four days of training on 888 A100 GPUs with a batch size of 646464. The hyper parameters for LoRA finetuning can be referenced as $rank=128$, with a learning rate of 1e-4. The learning rate for training the action control module is 1e-5. We only apply classifier free guidance [18] to the text prompt conditional input and use DDIM [39] sampling with 505050 sampling steps.
[…]
@@ 7 Conclusion
Acknowledgements
This work is partially supported by the National Nature Science Foundation of China (No. 62402406).
@@ A.2 Data PreProcessing
After obtaining the full video, we preprocess it in two steps: (1) Slicing: Randomly sample consecutive sequences of $k$ frames from the full video to form new video clips. Sampling continues until the total number of sampled frames reaches $n$ times the total frame count of the original video. In our experimental setup, $k=81$, which is slightly larger than the generation length of the video diffusion model (777777), and $n=3$; (2) Text Annotation: We use the open-source multimodal large language model MiniCPM-V [51] to annotate the sliced video clips with text. The corresponding prompts and annotation examples are shown in Fig. 11.

### v4

@@ 3 Preliminaries
We adopt a transformer-based latent video diffusion model [33, 31] as the backbone. Let $\mathbf{X}$ represent a video sequence. To reduce modeling complexity, an Encoder $E(\cdot)$ compresses the video spatially and temporally into a latent representation $\mathbf{Z}=E(\mathbf{X})$. With temporal compression ratio $r$, a $(1+rn)$-frame video is compressed into $(1+n)$ latent frames. Denoting the $i$-th frame as $\mathbf{x}^{i}$ and $i$-th latent as $\mathbf{z}^{i}$, we have $\mathbf{X}=[\mathbf{x}^{0},\mathbf{x}^{1},...,\mathbf{x}^{rn}]$ and $\mathbf{Z}=[\mathbf{z}^{0},\mathbf{z}^{1},...,\mathbf{z}^{n}]$. During training, noise is added to the clean latent $\mathbf{Z}_{0}$ to get noisy latent $\mathbf{Z}_{t}$ at timestep $t$. $\boldsymbol{\epsilon}$ is the added random noise. When considering action control, the action $\mathbf{A}=[\mathbf{a}^{1},\mathbf{a}^{2},...,\mathbf{a}^{rn}]$, where $\mathbf{a}^{i}$ represents the action taken at timestep (i1)(i-1) to transfer from $\mathbf{x}^{i-1}$ to $\mathbf{x}^{i}$. The corresponding action-conditioned loss function is:
@@ 4.2 Action Control Module
We incorporate action control modules into the transformer blocks of the video diffusion model to enable action-controllable generation. The architecture of the transformer block is shown in Fig. 3 (a), and the structure of the action control module is shown in Fig. 3 (b). Suppose that the input action includes both continuous mouse movement action $\mathbf{M}\in\mathbb{R}^{rn\times d_{1}}$, and discrete keyboard action $\mathbf{K}\in\mathbb{R}^{rn\times d_{2}}$. The intermediate feature in the Transformer is denoted as $\mathbf{F}\in\mathbb{R}^{(n+1)\times l\times c}$. The total number of video frames is $(1+rn)$, compressed to $(1+n)$ latent frames, where $r$ denotes the temporal compression ratio. $d_{1}$ and $d_{2}$ denote the dimension numbers of actions, $l$ is the length of token sequence and $c$ is the number of feature channels.
❑ Grouping Actions with a Sliding Window. Due to the temporal compression ratio $r$, the number of actions ($rn$) differs from the number of features ($n+1$), creating a granularity mismatch for action-feature fusion. As shown in Fig. 4, we address this by grouping actions using a sliding window of size $w$. For the $i$-th feature $\mathbf{f}^{i}$, we consider actions within $\mathbf{a}^{r\times(i-w+1)},...,\mathbf{a}^{ri}]$. This window design captures delayed action effects, such as how a jump command influences multiple subsequent frames. For out-of-range indices, boundary actions are used as padding. For mouse movement $\mathbf{M}$, the grouped action is $\mathbf{M}_{group}\in\mathbb{R}^{(n+1)\times rw\times d_{1}}$. As for keyboard actions $\mathbf{K}$, we first learn the embedding of actions and add the positional encoding. After that, we perform a grouping operation on the action embeddings to get $\mathbf{K}_{group}\in\mathbb{R}^{(n+1)\times rw\times c}$.
@@ 4.3 Autoregressive Generation of Long Action-Controllable Game Videos
Figure 5: Illustration of autoregressive video generation. The frames from index 0 to $k$ serve as conditional frames, while the remaining NkN-k frames are for prediction, with $k$ randomly selected. (a) Training stage: Loss computation and optimization focus only on the noise of predicted frames. (b) Inference stage: The model iteratively selects the latest $k+1$ frames as conditions to generate NkN-k new frames, enabling autoregressive generation.
@@
As shown in Fig. 5 (a), during training with $N+1$ frame latents (from 0 to $N$), we randomly select the first $k+1$ frames as conditions without adding noise, while adding noise only to the remaining NkN-k frames for noise prediction training. Although the first $k+1$ frames are input to the model, since they are assumed to be already generated, their predicted noise outputs are not utilized. To improve training efficiency, we focus on computing training losses only for the NkN-k frames that require prediction. As for inference, as shown in Fig. 5 (b), after full-sequence generation of the first $N+1$ frame latents, we can autoregressively generate new NkN-k frames by selecting the most recent $k+1$ frames from history video latents as conditions for each subsequent generation, and merge them into the history video latents. This process can be repeated to achieve infinite-length video generation. Unlike conventional next-frame generation methods [8, 42, 11], our approach supports multi-frame generation in one step, greatly reducing the time for long video generation.
@@ 5.2 Multi-Phase Training Strategy
Resolution $280\times 150$ 240240p 720720p $640\times 360$ 720720p 720720p $640\times 360$
@@ 6.1 Implementation Details
❑ Pretrained Model Setting. Our experiments are based on an internal 11B-sized transformer-based text-to-video diffusion model for research purpose, which is distilled from a larger pretrained video diffusion model and possesses strong generative priors in the open domain. The resolution of the game videos is $360\times 640$. The temporal compression rate of the VAE is $r=4$.
❑ Training and Inference Setting. Each phase of fine-tuning or training requires about two-four days of training on $8$ A100 GPUs with a batch size of $64$. The hyper parameters for LoRA finetuning can be referenced as $rank=128$, with a learning rate of 1e-4. The learning rate for training the action control module is 1e-5. We only apply classifier free guidance [18] to the text prompt conditional input and use DDIM [39] sampling with $50$ sampling steps.
❑ Evaluation Setting. We retained $5\%$ of our collected, segmented dataset as a testset, excluding it from training, and selected three subsets for ablation study: (1) only-key: contains only keyboard actions, designed to test the model’s ability to follow discrete actions; (2) mouse-small: includes small-scale continuous mouse movements; and (3) mouse-large: includes large-scale continuous mouse movements. Additionally, for qualitative experiments, we support custom combinations of input actions, allowing us to test complex or rare action combinations. We use these evaluation metrics: (1) Flow: calculates the optical flow of the generated video to reflect its dynamics, assessing action-following ability by measuring mean square error to the optical flow of the reference video; (2) Cam: computes the Euclidean distance between camera poses extracted from predicted videos and those extracted from reference videos, where both sets of camera poses are obtained using GLOMAP [32]. (3) CLIP: computes feature similarity in the CLIP [37] space to evaluate semantic relevance to the given text prompt; (4) FID, FVD: measures distribution differences between the generated videos and the reference, providing an assessment of generation quality.
@@ A.2 Data PreProcessing
After obtaining the full video, we preprocess it in two steps: (1) Slicing: Randomly sample consecutive sequences of $k$ frames from the full video to form new video clips. Sampling continues until the total number of sampled frames reaches $n$ times the total frame count of the original video. In our experimental setup, $k=81$, which is slightly larger than the generation length of the video diffusion model ($77$), and $n=3$; (2) Text Annotation: We use the open-source multimodal large language model MiniCPM-V [51] to annotate the sliced video clips with text. The corresponding prompts and annotation examples are shown in Fig. 11.